OpenStudy (anonymous):

3.34 g of an unknown acid is dissolved in water to make 250 mL solution. While performing the titration, 46.43 mL of the acid is needed to neutralize 7.41 mL of 0.50 M NaOH is a titration. What is the molar mass of the dry acid? (Assume the acid is monoprotic)

OpenStudy (aaronq):

You first need to find the molarity of the acid. Knowing the moles you can trace back to the molar mass. You need these two equations: \(M=\dfrac{n_{solute}}{L_{solution}}\) and \(n=\dfrac{m}{MW}\)

OpenStudy (anonymous):

Is the answer 167 g/mol? @aaronq

OpenStudy (aaronq):

i dont know, i havent worked it out. you're welcome to post your work if you'd like me to check it.

OpenStudy (anonymous):

Well using your hint. And some of the things that my teach taught me. I did: \[M_{a}=?\]\[V_{a}=46.43 mL\]\[M_{b}=0.50M\]\[V_{b}=7.41mL\] \[M_{a}*(46.43mL)=(0.50M)*(7.41mL)\] \[M_{a}=0.080 M\] \[\frac{ 0.080mol }{ 1.00L }=\frac{ 3.34g }{ 0.25L }* Xmol/g\]\[Xg/mol=\frac{ 3.34g*1.00L }{ 0.25L*0.080mol }\] The molar mass of the unknown acid is 167 g/mol. @aaronq

OpenStudy (anonymous):

teacher*

OpenStudy (aaronq):

\(\checkmark\) that is correct, nice formatting too.

OpenStudy (anonymous):

:D Thank you for the help @aaronq

OpenStudy (aaronq):

no problem dude!